# The Reliability of Remotely Piloted Aircraft System Performance under Communication Loss and Latency Uncertainties

### v1

@@ Abstract
Mission-critical use of highly maneuverable Remotely Piloted Aircraft Systems (RPAS) requires a thorough understanding of the reliability of their communication systems. Investigations into system-level performance under stochastic aviation communication conditions are critical for estimating mission success rates and assessing the risks associated with integrating RPAS into existing airspace, ensuring overall aviation safety. This study aims to quantify the impact of communication latency and complete signal loss on the mission completion performance of a highly maneuverable RPAS. The mission is defined as a static waypoint tracking task in three-dimensional airspace. We start with examining and deriving mathematical formulations of key reliability metrics of Required Communication Performance (RCP). These stochastic factors are then embedded into flight control simulations (i.e., communication availability and latency) to examine the system behavior. Lastly, we generate mission success rate and mission completion time envelopes through extensive multiprocessing Monte Carlo simulations through high-performance computing. We discover a drastic deterioration in flight performance while latency or availability erodes the stability margin. In addition, we propose a new reliability metric, namely communicability, which integrates three key RCP metrics and helps understanding the maximum tolerable latency to flight control. The procedure and results obtained from this research inform engineers designing RPAS with better trade-off between communication capability and flight control performance. Future works includes exploring alternative flight simulators (i.e., nonlinear dynamic inversion) with other missions (i.e., dynamic waypoint following), or develop delay-compensated optimal controls. The analysis on stability margin is also desired for theoretical verification. The code used in this research can be found from this LINK.
@@ 1 Introduction
The Remotely Piloted Aircraft System (RPAS) stands for the end-to-end system that an unmanned Remotely Piloted Aircraft (RPA) is piloted from a Remote Pilot Station (RPS) by a Remote Pilot (RP). Such systems can carry cameras, sensors, communications equipment, or other payloads to perform specific tasks. Although RPAS have been used for reconnaissance and information gathering of military operations since the 1950s (i.e., High-Altitude Long Endurance unmanned aerial vehicle) (Ma’Sum et al., 2013; Chaturvedi et al., 2019), their future role has expanded considerably to the integration of RPAS into existing civilian airspace due to the gradual maturity of aerospace autonomy capabilities (Kephart and Braasch, 2010). The growing interest in RPAS stems from the versatility and operational advantages. RPAS can undertake missions in hazardous environments without placing aircrews at physical risk, thus reducing the threat to fatalities. They are also invaluable in tasks such as real-time surveillance (Iscold et al., 2010; Thammawichai et al., 2017), search and rescue (Scherer et al., 2015), and environmental monitoring (Asadzadeh et al., 2022). The ability of RPAS to remain airborne for extended periods while providing constant updates and imagery makes them an increasingly important tool for both the official and the private sector (Abeyratne and Abeyratne, 2019).
@@
Air Traffic Management (ATM) serves as the backbone of safe and efficient flight operations within the active field of research on RPAS. ATM consists of three major functions (air traffic services, airspace management, and air traffic flow management) to ensure safe and efficient travel of aircraft within the given section of airspace (ICAO, 2006), by integrating human, software and hardware supported by communications, navigation, surveillance (CNS) capabilities. Air traffic services provides services to support the daily aeronautical operations such as collision avoidance, conflict resolution, and maintain aeronautical separation minima via timely information provision (ICAO AMCP WG-C1/WP8, 2011). In recent years, emerging aviation concept such as Advanced Air Mobility (AAM), is pushing the boundaries of ATM adaptation even further. Research institutes and government agencies such as the Federal Aviation Administration (FAA), National Aeronautics and Space Administration (NASA), and the European Union Aviation Safety Agency (EASA) are leading the research effort on Advanced Air Mobility (AAM) (Goyal and Cohen, 2022; Johnson and Silva, 2022). AAM pushes air transportation to the new stage by integrating physical and digital infrastructures to efficiently move the people and cargo in the airspace, with multi-rotor vertical or short takeoff and landing aircraft ideally being electrical (eVTOL/STOL). Urban Air Mobility (UAM) is viewed as the subset of AAM where the operation are mostly at urban scenarios (MITRE Engenuity, 2023). The Urban Mobility Concept of Operation defines the AAM operates above the AAM corridors at 400 feet above ground level (AGL), and cruises between $1,500$ and $4,000$ AGL at the speed between 50 knots and 140 knots. Besides normal data and voice communication requirements in ATM, video streams communication is the desired function for, (a) real-time monitoring of passenger cabins; (b) command and control (C2) link of pilot video; (c) passenger entertainments (FAA, 2023). To address the increased demand for AAM/UAM operations in the urban scenario, the communication system that can works smartly becomes vital to maintain a proper workload for the controller maintaining safety and efficiency in the performance-based airspace.
@@
However, this expanding role for RPAS comes with significant challenges to the community, particularly regarding the safe and seamless integration into the existing Air Traffic Management (ATM) system. One needs to ensure that RPAS can safely share the airspace without causing disruptions to manned aircraft. This requires the establishment of airworthiness standards and verification techniques that do not impose undue burdens on air traffic services. The lack of the standard regulations of integrating remotely piloted aircraft systems (RPAS) into existing airspace has been identified as a major challenge (Organization, 2017; Serafino et al., 2019). The integration of RPAS has been outline by the Federal Aviation Administration (FAA) (Huerta, 2013) and Single European Sky ATM Research (SESAR) joint undertaking (SESAR-JU, 2019), as well as the International Civil Aviation Organization (ICAO). The long-term objective is to ensure all RPAS are regulated as routine civil operations (Teutsch and Petersen, 2022). Under this big picture, the safety guarantee of RPAS system reliability remains a challenge, along with detect and avoid system verification, and human factor satisfactory. The command and control (C2) link performance on communication latency, availability, and continuity can be evaluated in real-time, while latency and availability are reported to be sufficient for communication link performance evaluation (Serafino et al., 2019).
@@
Researchers have been working on studying the reliability of safety-critical sub-systems in aviation (Pettit et al., 2001; Weckman et al., 2001; Lu et al., 2015; Muecklich et al., 2023), which also extends to the emerging field of system-level verification in either deterministic or probabilistic sense (Tomlin et al., 2001; Reinhardt, 2013; Hook et al., 2016). Aircraft performance is one of the prime examples for verification of reliability, as they can be modeled in the simulation environment (Heidlauf et al., 2018). However, existing works usually adopt simple approximations which assume the aircraft can be modeled with a point mass (Hu et al., 2020; Guo et al., 2021; Hu et al., 2022) or reduced number of freedom (i.e., Dubins Model) (Dubins, 1957; Molnar et al., 2024). On the other hand, communication latency and complete signal loss are critical events that impact RPAS system reliability, and drastically deteriorate flight performance, especially during autonomous operations. Additionally, the desired level of automation during certain RPAS missions is also an open-ended question to the community. For instance, three levels of control architectures are in service, (i) human-in-the-loop (HITL), where the control system requires continuous and direct human input; (ii) human-on-the-loop (HOTL), where human act as the supervisor to the machine; (iii) fully autonomous, where the RPAS determines procedures to fulfill mission requirements. RPAS are moving objects, proper autonomous functions needs to be put onboard to maintain stability as well as perform basic maneuvers. Studying the impact of C2 link issues helps answer the question of, how much autonomy and what type of control architecture do we need to be have onboard for RPAS in order to complete the given task?
@@ 2.1 Reliability of Communication
The special committee on Future Air Navigation System (FANS) of ICAO has identified the enhancement of communications, navigation, surveillance and air traffic management (CNS/ATM) as one of the first objectives (ICAO AMCP WG-C1/WP8, 2011). C2 link relies on signal propagation to transmit either higher-level operational instructions or immediate flight control parameters for Air-to-Ground or Air-to-Air communications. Signal strength estimation is one of the most critical part of signal propagation (Alsamhi et al., 2018) to ensure the receiver gaining sufficient level of power. Signal propagation can be affected by three main phenomena: reflection, diffraction, and scattering. Reflection occurs when a signal bounces off large, smooth surfaces, causing multipath fading due to multiple signal paths arriving at different times. Diffraction involves the bending of signals around sharp edges, allowing them to reach shadowed areas not in direct line of sight but causing attenuation in the process. Scattering happens when signals encounter small or irregular objects, dispersing energy in multiple directions and reducing signal strength, yet sometimes assisting in overcoming obstacles by reaching otherwise blocked areas.
@@
Specifically, the term communication stands for accurate information transfer between the transmitter and receiver. In this work, voice and data communications are considered within the scope of ATM. Voice communication for Air-to-Ground (A2G) systems within the continental U.S. airspace are two-way between the pilots and the ground stations, achieved by Remote Communication Facilities (RCFs) maintained by FAA. RCFs use Very High Frequency (VHF) and Ultra High Frequency (UHF) antenna transmitters and receivers mounted on buildings or towers. VHF operates from 118 MHZ to 136.975 MHZ are for civil aviation use, while UHF channel takes effect from 225 MHZ to 400 MHZ of military usage (Markus and Eftekari, 2020). High Frequency (HF) aeronautical voice communication also serves A/G communications for a much long range. It is available over international airspace (or continental U.S. only in the case of emergency), which covers frequency from 3 MHZ to 30 MHZ. Iridium and Inmarsat provide complete or approximate world-wide satellite voice communication (SATVOICE) service in Lower Earth Orbit (LEO) or Geostationary Earth Orbit (GEO). On the other side, data communication refers to textual data providing routes information to the pilots, which also referred as data link. Similar to voice communication, there is HFDL/VDL refers to data link using HF/VHF, as well as Iridium and Inmarsat providing satellite-based data communication (SATCOM) services primarily used in oceanic and remote airspace (The Performance Based Operations Aviation Rulemaking Committee (2018), PARC).
@@
The concept, Reliability, stands for the probability that the system functioning correctly under certain operation conditions over a specific period of time (Villemeur, 1992; Ahmad et al., 2017). The reliability of aviation communication system has been investigated by either the RCP concept by ICAO for manned commercial aviation, or the Required Link Performance (RLP) by the Joint Authorities for Rulemaking of Unmanned Systems (JARUS). RCP and RLP identical key metrics to quantify the reliability of communication systems, transaction time, availability, continuity, and integrity. Similar metrics has also been used in measuring reliability of satellite-based Differential Global Positioning System (DGPS) for maritime navigation (Moore et al., 2002) and other navigation application domains. The detailed explanations and mathematical derivations is given in Section 3, along with our newly proposed metric, communicability, to consider the first three key metrics together. For consistency of terminologies, we use RCP to represent RLP throughout this paper.
@@ 2.2 Communication Latency
Possible solutions to fight latencies are improving robot design, employing predictive modeling, and advancing communication technologies. For instance, the typical approach involves inserting a predicted compensatory delay into the feedback loop of the control system, allowing the operator to experience instantaneous synthetic feedback, which is particularly effective when latency is stable and predictable (Zhao et al., 2015; Wang et al., 2017). Emerging high-frequency communication technologies (i.e., 5G), promise both high data throughput and ultra-low latency, enabling more responsive and precise teleoperations (Voigtländer et al., 2017). Additionally, control augmentation strategies such as trajectory prediction (Pang et al., 2021, 2022) can alleviate latency impacts. However, unpredictable external disturbances can make these predictive models unreliable (Chen et al., 2007). Long-term solutions point toward autonomous operations, where vehicles sense, decide, and maneuver without human interference (Byrnes, 2014; Mayer, 2015). A hybrid approach that combines the aforementioned solution strategies may offer promising avenues for mitigating latency, with further necessary improvements.
@@ 2.3 Communication Loss
Many factors can cause complete communication signal loss for RPAS, which requires accurate measurement of sensor input (i.e., radar altimeters) for awareness of situations within the area of interest. The sensor data can be inaccurate, delayed, corrupted, or spoofed. Cyber-physical attacks toward communication infrastructures and links is the major source of complete communication signal loss (Dave et al., 2022; Ukwandu et al., 2022). Based on the awareness of the attacker blocking communication channels, we classify them into intentional and unintentional attacks. Intentional attacks toward aviation communication systems include jamming, spoofing, and tampering. Signal jamming refers to deliberately disrupting radio signals using devices that broadcast noise or create interference on the same frequency as aviation communication channels. Spoofing injects false signals to mimic legitimate communication channels to disrupt ATC and pilots. Tempering, also known as man-in-the-middle attack, refers to unauthorized access to the system or even damage physical communication infrastructures. Distributed-Denial-of-Service (DDoS) is viewed as a special form of spoofing by massively injecting false signals into the communication system. Unintentional attacks include the electromagnetic interference (EMI) caused by onboard equipment or passenger electronics, cross-modulation caused the mixing of signals from different frequencies. Space weather events is one example of unintentional attacks towards the aviation communication systems. Space weather events disrupt ionosphere reflections and create blockage of shortwave radio transmissions (Xue et al., 2024). The formation of space weather events is highly unpredictable, as it is related to the moving of particles inside of the Earth’s magnetic field towards the poles caused by solar activities (Tsurutani et al., 2005, 2009, 2022). North Polar Flights traversing the area north of $82^{o}$ primarily rely on HF to maintain connection with ground stations, or SATCOM for most remote areas. However, both HF and SATCOM are heavily impacted by space weather events in the area north of $82^{o}$, and even inaccessible for HF bands below 30MHz (Xue et al., 2024). In these scenarios, flights are required to divert or land near the North Pole due to the lack of timely and accurate space weather event prediction.
@@
3 Required Navigation Performance (RCP)

### v2

# The Reliability of Remotely Piloted Aircraft System Performance under Aeronautical Communication Uncertainties

@@ Abstract
Mission-critical operations of highly maneuverable Remotely Piloted Aircraft Systems (RPAS) require reliable communication to ensure safe integration into existing airspace. Understanding system-level performance under stochastic communication conditions is essential for estimating mission success and assessing operational risks. This study quantifies the impact of communication latency and complete signal loss on the mission completion performance of a highly maneuverable RPAS. The mission is defined as a static waypoint tracking task in three-dimensional airspace. We first derive mathematical formulations for key reliability metrics within the Required Communication Performance (RCP) framework. These stochastic communication factors, including latency and availability, are then incorporated into flight control simulations to evaluate system behavior. Extensive multiprocessing Monte Carlo simulations are conducted using high-performance computing to generate mission success rate and mission completion time envelopes. Results show significant degradation in flight performance as communication latency increases or availability decreases, which directly reduces the system stability margin. To better characterize this relationship, we introduce a new reliability metric, communicability, which integrates three key RCP metrics and provides insight into the maximum tolerable latency for flight control. The proposed framework informs RPAS design by revealing trade-offs between communication capability and flight control performance. The code used in this study is publicly available at this repository.
@@ 1 Introduction
The Remotely Piloted Aircraft System (RPAS) is an end-to-end system in which an unmanned Remotely Piloted Aircraft (RPA) is flown from a Remote Pilot Station (RPS) by a Remote Pilot (RP). An RPAS can carry cameras, sensors, communication equipment, or other payloads to perform specific tasks. Although RPAS have been used for reconnaissance and information gathering of military operations since the 1950s (i.e., High-Altitude Long Endurance unmanned aerial vehicle) (Ma’Sum et al., 2013; Chaturvedi et al., 2019), their future role has expanded considerably to the integration of RPAS into existing civilian airspace due to the gradual maturity of aerospace autonomy capabilities (Kephart and Braasch, 2010). The growing interest in RPAS stems from the versatility and operational advantages. RPAS can undertake missions in hazardous environments without placing aircrews at physical risk, thus reducing the threat to fatalities. They are also invaluable in tasks such as real-time surveillance (Iscold et al., 2010; Thammawichai et al., 2017), search and rescue (Scherer et al., 2015), and environmental monitoring (Asadzadeh et al., 2022). The ability of RPAS to remain airborne for extended periods while providing constant updates and imagery makes them an increasingly important tool for both the official and the private sector (Abeyratne and Abeyratne, 2019).
@@
Air Traffic Management (ATM) serves as the backbone of safe and efficient flight operations within the active field of research on RPAS. ATM consists of three major functions (air traffic services, airspace management, and air traffic flow management) to ensure safe and efficient travel of aircraft within the given section of airspace (ICAO, 2006), by integrating human, software and hardware supported by communications, navigation, surveillance (CNS) capabilities (Flavio Vismari and Camargo Junior, 2011). Air traffic services provides services to support the daily aeronautical operations such as collision avoidance, conflict resolution, and maintain aeronautical separation minima via timely information provision (ICAO AMCP WG-C1/WP8, 2011). In recent years, emerging aviation concept such as Advanced Air Mobility (AAM), is pushing the boundaries of ATM adaptation even further. Research institutes and government agencies such as the Federal Aviation Administration (FAA), National Aeronautics and Space Administration (NASA), and the European Union Aviation Safety Agency (EASA) are leading the research effort on Advanced Air Mobility (AAM) (Goyal and Cohen, 2022; Johnson and Silva, 2022). AAM pushes air transportation to the new stage by integrating physical and digital infrastructures to efficiently move the people and cargo in the airspace, with multi-rotor vertical or short takeoff and landing aircraft ideally being electrical (eVTOL/STOL). Urban Air Mobility (UAM) is viewed as the subset of AAM where the operation are mostly at urban scenarios (MITRE Engenuity, 2023). The Urban Mobility Concept of Operation defines the AAM operates above the AAM corridors at 400 feet above ground level (AGL), and cruises between $1,500$ and $4,000$ AGL at the speed between 50 knots and 140 knots. Besides normal data and voice communication requirements in ATM, video streams communication is the desired function for, (a) real-time monitoring of passenger cabins; (b) command and control (C2) link of pilot video; (c) passenger entertainments (FAA, 2023). To address the increased demand for AAM/UAM operations in the urban scenario, the communication system that can works smartly becomes vital to maintain a proper workload for the controller maintaining safety and efficiency in the performance-based airspace.
@@
However, this expanding role for RPAS comes with significant challenges to the community, particularly regarding the safe and seamless integration into the existing ATM system. One needs to ensure that RPAS can safely share the airspace without causing disruptions to manned aircraft. This requires the establishment of airworthiness standards and verification techniques that do not impose undue burdens on air traffic services. The lack of the standard regulations of integrating remotely piloted aircraft systems (RPAS) into existing airspace has been identified as a major challenge (Organization, 2017; Serafino et al., 2019). The integration of RPAS has been outlined by the Federal Aviation Administration (FAA) (Huerta, 2013) and Single European Sky ATM Research (SESAR) joint undertaking (SESAR-JU, 2019), as well as the International Civil Aviation Organization (ICAO). The long-term objective is to ensure all RPAS are regulated as routine civil operations (Teutsch and Petersen, 2022). Under this big picture, the safety guarantee of RPAS system reliability remains a challenge, along with detect and avoid system verification, and human factor satisfactory. The command and control (C2) link performance on communication latency, availability, and continuity can be evaluated in real-time, while latency and availability are reported to be sufficient for communication link performance evaluation (Serafino et al., 2019). Safety-critical communication networks have been the subject of reliability modeling in other domains, including nuclear power plant instrumentation and control systems (Lee et al., 2015) and fieldbus-based distributed control systems (Cauffriez et al., 2004), providing transferable insights for RPAS communication dependability.
@@
Researchers have been working on studying the reliability of safety-critical sub-systems in aviation (Pettit et al., 2001; Weckman et al., 2001; Lu et al., 2015; Muecklich et al., 2023). Complementary efforts in the reliability community have advanced the understanding of accident causation and system safety through both organizational and technical lenses (Saleh et al., 2010), analyzed trends in aviation maintenance risk (Marais and Robichaud, 2012), and developed systematic approaches to model general aviation accidents such as high-risk occurrence chains in helicopter operations (Rao and Marais, 2018) and state-based accident models (Rao and Marais, 2020). These lines of research also extend to the emerging field of system-level verification in either deterministic or probabilistic sense (Tomlin et al., 2001; Reinhardt, 2013; Hook et al., 2016). Aircraft performance is one of the prime examples for verification of reliability, as they can be modeled in the simulation environment (Heidlauf et al., 2018). However, existing works usually adopt simple approximations which assume the aircraft can be modeled with a point mass (Hu et al., 2020; Guo et al., 2021a; Hu et al., 2022) or reduced number of freedom (i.e., Dubins Model) (Dubins, 1957; Molnar et al., 2024). On the other hand, communication latency and complete signal loss are critical events that impact RPAS system reliability, and drastically deteriorate flight performance, especially during autonomous operations. Additionally, the desired level of automation during certain RPAS missions is also an open-ended question to the community. For instance, three levels of control architectures are in service, (i) human-in-the-loop (HITL), where the control system requires continuous and direct human input; (ii) human-on-the-loop (HOTL), where human act as the supervisor to the machine; (iii) fully autonomous, where the RPAS determines procedures to fulfill mission requirements. RPAS are moving objects, proper autonomous functions needs to be put onboard to maintain stability as well as perform basic maneuvers. Studying the impact of C2 link issues helps answer the question of, how much autonomy and what type of control architecture do we need to be have onboard for RPAS in order to complete the given task?
@@ 2.1 Reliability of Aeronautical Communication
The special committee on Future Air Navigation System (FANS) of ICAO has identified the enhancement of communications, navigation, surveillance and air traffic management (CNS/ATM) as one of the first objectives (ICAO AMCP WG-C1/WP8, 2011). C2 link relies on signal propagation to transmit either higher-level operational instructions or immediate flight control parameters for Air-to-Ground (A2G) or Air-to-Air (A2A) communications. Signal strength estimation is one of the most critical part of signal propagation (Alsamhi et al., 2018) to ensure the receiver gaining sufficient level of power. Signal propagation can be affected by three main phenomena: reflection, diffraction, and scattering. Reflection occurs when a signal bounces off large, smooth surfaces, causing multipath fading due to multiple signal paths arriving at different times. Diffraction involves the bending of signals around sharp edges, allowing them to reach shadowed areas not in direct line of sight but causing attenuation in the process. Scattering happens when signals encounter small or irregular objects, dispersing energy in multiple directions and reducing signal strength, yet sometimes assisting in overcoming obstacles by reaching otherwise blocked areas.
@@
Specifically, the term communication stands for accurate information transfer between the transmitter and receiver. Voice and data communications are considered within the scope of current ATM services. Voice communication for A2G systems within the continental U.S. airspace are two-way between the pilots and the ground stations, achieved by Remote Communication Facilities (RCFs) maintained by FAA. RCFs use Very High Frequency (VHF) and Ultra High Frequency (UHF) antenna transmitters and receivers mounted on buildings or towers. VHF operates from 118 MHZ to 136.975 MHZ are for civil aviation use, while UHF channel takes effect from 225 MHZ to 400 MHZ of military usage (Markus and Eftekari, 2020). High Frequency (HF) aeronautical voice communication also serves A/G communications for a much long range. It is available over international airspace (or continental U.S. only in the case of emergency), which covers frequency from 3 MHZ to 30 MHZ. Iridium and Inmarsat provide complete or approximate world-wide satellite voice communication (SATVOICE) service in Lower Earth Orbit (LEO) or Geostationary Earth Orbit (GEO). The reliability of satellite communication systems has been analyzed using Markov Bayesian Networks to model subsystem interactions and optimize redundancy designs in the space environment (Chen et al., 2024). On the other side, data communication refers to textual data providing routes information to the pilots, which also referred as data link. Similar to voice communication, there is HFDL/VDL refers to data link using HF/VHF, as well as Iridium and Inmarsat providing satellite-based data communication (SATCOM) services primarily used in oceanic and remote airspace (The Performance Based Operations Aviation Rulemaking Committee (PARC), 2018).
@@
The concept, reliability, stands for the probability that the system functioning correctly under certain operation conditions over a specific period of time (Villemeur, 1992; Ahmad et al., 2017). The intellectual roots and evolution of reliability engineering as a discipline have been well documented (Saleh and Marais, 2006a), and the methods for system reliability optimization have advanced considerably over several decades (Coit and Zio, 2019). Understanding the financial implications of reliability is also important, as the net present value framework illustrates how system reliability directly impacts revenue generation capability and investment decisions (Saleh and Marais, 2006b). The reliability of aviation communication system has been investigated by either the RCP concept by ICAO for manned commercial aviation, or the Required Link Performance (RLP) by the Joint Authorities for Rulemaking of Unmanned Systems (JARUS). RCP and RLP have identical key metrics to quantify the reliability of communication systems as, transaction time, availability, continuity, and integrity. Similar metrics has also been used in measuring reliability of satellite-based Differential Global Positioning System (DGPS) for maritime navigation (Moore et al., 2002) and other navigation application domains. The detailed explanations and mathematical derivations is given in Section 3, along with our newly proposed metric, communicability, to consider the first three key metrics together. For consistency of terminologies, we use RCP instead of RLP throughout this paper.
@@ 2.2 Communication Latency
Possible solutions to fight latencies are improving robot design, employing predictive modeling, and advancing communication technologies. In related transportation domains, model-based verification methods have been developed to address parameter uncertainty in train control systems where communication delays and packet-loss factors directly affect safety (Cheng et al., 2016), and formal verification using Colored Petri Nets has been applied to validate safety communication protocols (Lijie et al., 2012). For instance, the typical approach involves inserting a predicted compensatory delay into the feedback loop of the control system, allowing the operator to experience instantaneous synthetic feedback, which is particularly effective when latency is stable and predictable (Zhao et al., 2015; Wang et al., 2017). Emerging high-frequency communication technologies (i.e., 5G), promise both high data throughput and ultra-low latency, enabling more responsive and precise teleoperations (Voigtländer et al., 2017). Additionally, control augmentation strategies such as trajectory prediction (Pang et al., 2021, 2022) can alleviate latency impacts. However, unpredictable external disturbances can make these predictive models unreliable (Chen et al., 2007). Long-term solutions point toward autonomous operations, where vehicles sense, decide, and maneuver without human interference (Byrnes, 2014; Mayer, 2015). A hybrid approach that combines the aforementioned solution strategies may offer promising avenues for mitigating latency, with further necessary improvements.
@@ 2.3 Communication Loss
Many factors can cause complete communication signal loss for RPAS, which requires accurate measurement of sensor input (i.e., radar altimeters) for awareness of situations within the area of interest. From an organizational perspective, communication loss events can be understood within the broader framework of risk dynamics, where operational pressures may erode safety margins over time (Marais and Saleh, 2008). Communication redundancy is a primary design countermeasure, though the value of such redundancy is contingent on the prevalence of common-cause failures that can simultaneously affect multiple communication channels (Hoepfer et al., 2009). The sensor data can be inaccurate, delayed, corrupted, or spoofed. Cyber-physical attacks toward communication infrastructures and links is the major source of complete communication signal loss (Dave et al., 2022; Ukwandu et al., 2022). Based on the awareness of the attacker blocking communication channels, we classify them into intentional and unintentional attacks. Intentional attacks toward aviation communication systems include jamming, spoofing, and tampering. Signal jamming refers to deliberately disrupting radio signals using devices that broadcast noise or create interference on the same frequency as aviation communication channels. Spoofing injects false signals to mimic legitimate communication channels to disrupt ATC and pilots. Tempering, also known as man-in-the-middle attack, refers to unauthorized access to the system or even damage physical communication infrastructures. Distributed-Denial-of-Service (DDoS) is viewed as a special form of spoofing by massively injecting false signals into the communication system. Unintentional attacks include the electromagnetic interference (EMI) caused by onboard equipment or passenger electronics, cross-modulation caused the mixing of signals from different frequencies. Space weather events is one example of unintentional attacks towards the aviation communication systems. Space weather events disrupt ionosphere reflections and create blockage of shortwave radio transmissions (Xue et al., 2024). The formation of space weather events is highly unpredictable, as it is related to the moving of particles inside of the Earth’s magnetic field towards the poles caused by solar activities (Tsurutani et al., 2005, 2009, 2022). North Polar Flights traversing the area north of $82^{o}$N primarily rely on HF to maintain connection with ground stations, or SATCOM for most remote areas. However, both HF and SATCOM are heavily impacted by space weather events in the area north of $82^{o}$N, and even inaccessible for HF bands below 30MHz (Xue et al., 2024). In these scenarios, flights are required to divert or land near the North Pole due to the lack of timely and accurate space weather event prediction.
@@
3 Required Communication Performance (RCP)